OpenStudy (anonymous):

how do i factor each polynomial in 30b^2+48b-2?

OpenStudy (jhannybean):

What is the common multiple in 30, 48 and 2?

OpenStudy (jhannybean):

hint: \(\large \color{green}{30}b^2+\color{green}{48}b-\color{green}2\)

OpenStudy (anonymous):

2

OpenStudy (jhannybean):

goood job :D

OpenStudy (jhannybean):

Is that what you were looking for?

OpenStudy (anonymous):

what do i do next?

OpenStudy (jhannybean):

so now you'll factor out a two and see what multiplies to 2 to get your original equation, \[\large 2(?b^2+?b-1)\]

OpenStudy (anonymous):

2(15b^2+24b-1)

OpenStudy (anonymous):

?

OpenStudy (jhannybean):

yes, good job. i forgot to mention that instead of factoring out a b, we could have divided everything by 2 also... So let's also try that, what would you do if you had \[\large \frac{30b^2}{2}+\frac{48b}{2}-\frac{2}{2}\] (it's just easier to solve i guess)

OpenStudy (anonymous):

what about w^5+4w^4+10w^3+40w^2

OpenStudy (jhannybean):

what is the GCF here? (the w with the smallest power)?

OpenStudy (anonymous):

i don't know

OpenStudy (jhannybean):

hmm... \[\large \color{green}{w^5}+4\color{green}{w^4}+10\color{green}{w^3}+40\color{green}{w^2}\] out of these which w has the smallest exponent?

OpenStudy (anonymous):

w2

OpenStudy (jhannybean):

good job, how you can divide everything by w^2

OpenStudy (anonymous):

i dont know

OpenStudy (jhannybean):

\[\large \frac{w^5}{w^2}+\frac{4w^4}{w^2}+\frac{10w^3}{w^2}+\frac{40w^2}{w^2}\]

OpenStudy (anonymous):

is that all i do?